# Benchmarking Quantum Data Center Architectures: A Performance and Scalability Perspective

### v1

@@ Abstract.
In this paper, we present a systematic benchmarking study of four representative QDC architectures—QFly, BCube, Clos, and Fat-Tree—quantifying their impact on distributed quantum circuit execution latency, resource contention, and scalability. Focusing on quantum-specific effects absent from classical data-center evaluations, we analyze how optical-loss-induced Einstein–Podolsky–Rosen (EPR) pair generation delays, coherence-limited entanglement retry windows, and contention from teleportation-based non-local gates shape end-to-end execution performance. Across diverse circuit workloads, we evaluate how architectural properties such as path diversity and path length, and shared BSM (Bell State Measurement) resources indteract with optical-switch insertion loss and reconfiguration delay. Our results show that distributed quantum performance is jointly shaped by topology, scheduling policies, and physical-layer parameters, and that these factors interact in nontrivial ways. Together, these insights provide quantitative guidance for the design of scalable and high-performance quantum data-center architectures for DQC.
@@ 1. Introduction
Distributed Quantum Computing (DQC) has emerged as a promising approach for scaling quantum systems beyond the limits of single quantum processors by interconnecting multiple smaller Quantum Processing Units (QPUs) through quantum networks (sakuma2024optical; cirac1999distributed; zhang2025switchqnet; shapourian2025quantum). In DQC, circuits are partitioned across QPUs, and non-local operations are executed using entanglement-assisted primitives such as gate or state teleportation. Recent experimental demonstrations across multiple qubit platforms have validated the feasibility of such non-local operations (main2025distributed; sinclair2025fault), while compiler and orchestration frameworks based on lattice surgery and distributed surface codes for error correction have further advanced fault-tolerant distributed quantum computing (watkins2024high; horsman2012surface; keskin2025lattice; cuomo2023optimized; zhang2025switchqnet).
@@
To support DQC at scale, several quantum data-center (QDC) architectures have been proposed, many inspired by classical data-center topologies but adapted to quantum networking constraints. These architectures emphasize different design goals, such as minimizing path length (e.g., Qfly (sakuma2024optical)) or maximizing path diversity (e.g.,Fat-tree (shapourian2025quantum)). In the classical domain, such topologies have been extensively benchmarked using metrics including throughput, fault tolerance, and capital efficiency (popa2010cost; liu2013data; xia2016survey). However, a comparable, cross-architectural evaluation remains largely absent in the quantum setting.
@@
Figure 1 illustrates how quantum-specific constraints arise from the physical realization of quantum interconnects and shape how entanglement is established between QPUs. In server-centric architectures such as BCube, communication paths form repeater chains in which intermediate QPUs actively perform entanglement swapping, introducing and causing latency and noise to accumulate across multiple hops and making EPR generation sensitive to scheduling under finite coherence windows (pouryousef2024minimal; li2025optimising). In contrast, switch-centric architectures route entanglement through the optical switching fabric, where QPUs act only as endpoints and swapping is performed at BSM-capable switches along the path. In this figure, for both cases, we abstract each non-local operation as consuming a single virtual entanglement link between the communicating QPUs, which aggregates physical losses and operational imperfections.
In this work, we present a systematic benchmarking of four representative QDC architectures-QFly (sakuma2024optical), BCube (shapourian2025quantum), Clos (shapourian2025quantum), and Fat-Tree (shapourian2025quantum)–under diverse circuit workloads and physical-layer conditions. Our evaluation explicitly examines how architectural design choices affect (i) end-to-end circuit execution latency under realistic entanglement-generation models, (ii) the number of switch hops incurred by non-local gates, and (iii) contention for BSM resources. We study multiple architectural variants to isolate the trade-offs between path length, contention, and scalability. Workloads span a range of circuit widths and depths, connectivity patterns, and two-qubit gate densities, enabling us to assess how algorithmic structure interacts with architectural and physical constraints.
@@ 2. Data Center Network Designs
Distributed quantum computing performance depends on multiple factors, including circuit partitioning (kaur2025optimized; burt2024generalised; baker2020time), compiler transformations (e.g., circuit rewrites that adapt programs to architectural constraints and reduce non-local communication) (keskin2025lattice), and entanglement-generation and scheduling policies (monroe2014large; jiang2007distributed). Modeling all such interactions jointly is beyond the scope of this work. Instead, our objective is to establish a controlled architectural baseline that isolates the impact of network topology and shared photonic resources on distributed execution.
@@ 2.1. QDC Hardware Components and Entanglement Generation
In this work, we focus on a scatter–scatter (midpoint-interference) entanglement-generation protocol, which is representative of near-term photonic quantum networks and is widely adopted in experimental and architectural studies. Alternative protocols—such as memory–memory or memory–source links—are discussed extensively in the literature (e.g., (beukers2024remote)) and are left for future exploration.
@@
This distinction has important system-level consequences. In a switch-centric design, each multi-hop optical path requires only a single BSM operation for an end-to-end Bell pair, whereas in a server-centric design, a repeater chain with $N_{\mathrm{rep}}$ intermediate QPUs requires $N_{\mathrm{rep}}+1$ Bell-state measurements and corresponding storage intervals. These additional operations directly impact achievable entanglement rates, fidelity, and scheduling complexity, motivating the architectural comparisons studied in the remainder of this paper. This distinction parallels classical data-center networks, where switch-centric Clos fabrics treat servers as pure endpoints, while server-centric designs such as DCell and BCube involve servers directly in routing and forwarding decisions (guo2008dcell; guo2009bcube). Analogously, in quantum data centers, switch-centric architectures centralize entanglement operations in the fabric, whereas server-centric architectures allow QPUs to participate in entanglement distribution.
@@ 2.3. Fat-Tree Architecture
Fat-Tree is a canonical switch-centric data-center network originally proposed in the classical networking literature to provide full bisection bandwidth (i.e., sufficient aggregate interconnect capacity to support simultaneous full-rate communication between any equal partition of endpoints), predictable routing, and high path diversity at scale (leiserson1985fattree; alfares2008fattree). Its design philosophy is to realize high aggregate bandwidth using many moderate-radix switches arranged in a symmetric multi-tier hierarchy, rather than relying on expensive ultra-high-radix switches.
@@
This construction follows the canonical $k$-ary Fat-Tree topology of Al-Fares et al. (al2008scalable).
@@ 2.5. QFly Architecture
QFly is a flattened, low-diameter switch-centric architecture inspired by Dragonfly-style interconnects, which employ groups of high-radix routers and sparse global links to achieve low diameter while limiting long-cable cost (kim2008technology). In QFly, QPUs are grouped under a smaller number of high-radix optical switches, reducing hierarchy depth and the number of fabric elements compared to Fat-Tree and Clos.
@@ 2.6. BCube Architecture
We follow the original BCube construction of Guo et al. (guo2009bcube). An $(n,k_{\mathrm{bcube}})$-BCube consists of $k_{\mathrm{bcube}}+1$ switch layers, each containing $n^{k_{\mathrm{bcube}}}$ switches of radix $n$. Each QPU connects to exactly one switch in every layer, and the total system capacity scales as $n^{k_{\mathrm{bcube}}+1}$ QPUs as shown in figure 5.
@@ 2.7. Optical Loss and EPR-Generation Timing Model
corresponding to a Beneš-style non-blocking optical switch design (spanke1987n). Let $n_{\mathrm{sw}}$ denote the number of switch traversals along the path; then $L_{\mathrm{sw}}=n_{\mathrm{sw}}\,L_{\mathrm{sw}}(k)$.
@@
In server-centric architectures such as BCube, entanglement must be generated on multiple segments and combined through entanglement swapping at intermediate QPUs. Because partially generated entanglement must be stored while waiting for other segments to succeed, we impose a coherence-limited retry window $\tau_{\mathrm{cut}}$. If an end-to-end EPR pair has not been successfully formed via swapping within $\tau_{\mathrm{cut}}$, the attempt is discarded and restarted. We evaluate two repeater-chain scheduling protocols—sequential and parallel—introduced in (pouryousef2024minimal), which capture different trade-offs between control simplicity and latency. For these protocols, we estimate the raw end-to-end entanglement delivery rate $R_{\mathrm{raw}}$ and compute
@@ 3. Evaluation
System size is the number of QPUs, denoted by $N$. The total circuit width is set proportional to system scale by assigning each QPU a fixed local grid of data qubits. We evaluate multiple system scales by varying $N$ while keeping the per-QPU data-qubit capacity and the number of communication qubits per QPU fixed. Thus, increasing $N$ increases the total circuit width proportionally (more QPUs, each with the same local qubit grid). The resulting circuits are then partitioned using the same compilation and mapping pipeline and simulated under architecture-specific routing and scheduling policies (we explain this later). Circuits are partitioned using a lightweight Kernighan–Lin (KL) heuristic (KL_algorithm). For the monolithic baseline, we assume a single QPU that hosts all physical qubits required to handle the circuit, supports full all-to-all connectivity, and incurs no inter-QPU communication or entanglement-generation overhead. The circuit execution latency in this case is denoted by $T_{\mathrm{mono}}$.
@@ 3.1. Sanity Checks: Remote-Gate Timing Under Varying Parameters
For the Server-centric paradigm, we evaluate both sequential (S) and parallel (P) swap-scheduling protocols proposed in (pouryousef2024minimal), with a memory cutoff time $\tau_{\mathrm{cut}}$ that limits how long a partially generated entangled state may be stored if the required swaps have not yet been completed, after which the attempt is discarded. We report two derived, dimensionless metrics to summarize the behavior of the optical loss and entanglement-timing model.
@@ 3.3. Architecture Instantiations
Following the design space described in Section 2, we instantiate Clos topologies using an automated layout procedure that derives a feasible configuration for each target system size $N$. The procedure selects an even switch radix $k$, determines the number of top-of-rack (ToR) switches $T=k^{2}/4$, and assigns the number of QPUs per rack $R$ such that $T\times R\geq N$, while respecting practical constraints on rack fanout and ToR port budgets. This construction follows the standard folded-Clos model, where each ToR switch allocates $k/2$ ports to QPUs and $k/2$ ports to uplinks (clos1953study; greenberg2009vl2; singh2015jupiter).
@@ 3.4. Sensitivity of BCube to communication qubits coherence time
In this experiment, we fix the circuit and network configuration (long-range workload, fixed number of QPUs, communication qubits, and BSM allocation), set the cutoff time $\tau_{\mathrm{cut}}$ in the BCube network model, simulate distributed circuit execution, and record the resulting execution latency. For BCube, we use the parallel swap scheduling protocol proposed in (pouryousef2024minimal).

### v2

@@ Abstract.
In this paper, we present a systematic benchmarking study of four representative QDC architectures—QFly, BCube, Clos, and Fat-Tree—quantifying their impact on distributed quantum circuit execution latency, resource contention, and scalability. Focusing on quantum-specific effects absent from classical data-center evaluations, we analyze how optical-loss-induced Einstein–Podolsky–Rosen (EPR) pair generation delays, coherence-limited entanglement retry windows, and contention from teleportation-based non-local gates shape end-to-end execution performance. Across diverse circuit workloads, we evaluate how architectural properties such as path diversity and path length, and shared BSM (Bell State Measurement) resources interact with optical-switch insertion loss and reconfiguration delay. Our results show that distributed quantum performance is jointly shaped by topology, scheduling policies, and physical-layer parameters, and that these factors interact in nontrivial ways. Together, these insights provide quantitative guidance for the design of scalable and high-performance quantum data-center architectures for DQC.
@@ 1. Introduction
Distributed Quantum Computing (DQC) has emerged as a promising approach for scaling quantum systems beyond the limits of single quantum processors by interconnecting multiple smaller Quantum Processing Units (QPUs) through quantum networks (Sakuma et al., 2024; Cirac et al., 1999; Zhang et al., 2025; Shapourian et al., 2025). In DQC, circuits are partitioned across QPUs, and non-local operations are executed using entanglement-assisted primitives such as gate or state teleportation. Recent experimental demonstrations across multiple qubit platforms have validated the feasibility of such non-local operations (Main et al., 2025; Sinclair et al., 2025), while compiler and orchestration frameworks based on lattice surgery and distributed surface codes for error correction have further advanced fault-tolerant distributed quantum computing (Watkins et al., 2024; Horsman et al., 2012; Keskin et al., 2025; Cuomo et al., 2023; Zhang et al., 2025).
@@
To support DQC at scale, several quantum data-center (QDC) architectures have been proposed, many inspired by classical data-center topologies but adapted to quantum networking constraints. These architectures emphasize different design goals, such as minimizing path length (e.g., Qfly (Sakuma et al., 2024)) or maximizing path diversity (e.g.,Fat-tree (Shapourian et al., 2025)). In the classical domain, such topologies have been extensively benchmarked using metrics including throughput, fault tolerance, and capital efficiency (Popa et al., 2010; Liu et al., 2013; Xia et al., 2016). However, a comparable, cross-architectural evaluation remains largely absent in the quantum setting.
@@
Figure 1 illustrates how quantum-specific constraints arise from the physical realization of quantum interconnects and shape how entanglement is established between QPUs. In server-centric architectures such as BCube, communication paths form repeater chains in which intermediate QPUs actively perform entanglement swapping, introducing and causing latency and noise to accumulate across multiple hops and making EPR generation sensitive to scheduling under finite coherence windows (Pouryousef et al., 2024; Li et al., 2025). In contrast, switch-centric architectures route entanglement through the optical switching fabric, where QPUs act only as endpoints and swapping is performed at BSM-capable switches along the path. In this figure, for both cases, we abstract each non-local operation as consuming a single virtual entanglement link between the communicating QPUs, which aggregates physical losses and operational imperfections.
In this work, we present a systematic benchmarking of four representative QDC architectures-QFly (Sakuma et al., 2024), BCube (Shapourian et al., 2025), Clos (Shapourian et al., 2025), and Fat-Tree (Shapourian et al., 2025)–under diverse circuit workloads and physical-layer conditions. Our evaluation explicitly examines how architectural design choices affect (i) end-to-end circuit execution latency under realistic entanglement-generation models, (ii) the number of switch hops incurred by non-local gates, and (iii) contention for BSM resources. We study multiple architectural variants to isolate the trade-offs between path length, contention, and scalability. Workloads span a range of circuit widths and depths, connectivity patterns, and two-qubit gate densities, enabling us to assess how algorithmic structure interacts with architectural and physical constraints.
@@ 2. Data Center Network Designs
Distributed quantum computing performance depends on multiple factors, including circuit partitioning (Kaur et al., 2025; Burt et al., 2024; Baker et al., 2020), compiler transformations (e.g., circuit rewrites that adapt programs to architectural constraints and reduce non-local communication) (Keskin et al., 2025), and entanglement-generation and scheduling policies (Monroe et al., 2014; Jiang et al., 2007). Modeling all such interactions jointly is beyond the scope of this work. Instead, our objective is to establish a controlled architectural baseline that isolates the impact of network topology and shared photonic resources on distributed execution.
@@ 2.1. QDC Hardware Components and Entanglement Generation
In this work, we focus on a scatter–scatter (midpoint-interference) entanglement-generation protocol, which is representative of near-term photonic quantum networks and is widely adopted in experimental and architectural studies. Alternative protocols—such as memory–memory or memory–source links—are discussed extensively in the literature (e.g., (Beukers et al., 2024)) and are left for future exploration.
@@
This distinction has important system-level consequences. In a switch-centric design, each multi-hop optical path requires only a single BSM operation for an end-to-end Bell pair, whereas in a server-centric design, a repeater chain with $N_{\mathrm{rep}}$ intermediate QPUs requires $N_{\mathrm{rep}}+1$ Bell-state measurements and corresponding storage intervals. These additional operations directly impact achievable entanglement rates, fidelity, and scheduling complexity, motivating the architectural comparisons studied in the remainder of this paper. This distinction parallels classical data-center networks, where switch-centric Clos fabrics treat servers as pure endpoints, while server-centric designs such as DCell and BCube involve servers directly in routing and forwarding decisions (Guo et al., 2008, 2009). Analogously, in quantum data centers, switch-centric architectures centralize entanglement operations in the fabric, whereas server-centric architectures allow QPUs to participate in entanglement distribution.
@@ 2.3. Fat-Tree Architecture
Fat-Tree is a canonical switch-centric data-center network originally proposed in the classical networking literature to provide full bisection bandwidth (i.e., sufficient aggregate interconnect capacity to support simultaneous full-rate communication between any equal partition of endpoints), predictable routing, and high path diversity at scale (Leiserson, 1985; Al-Fares et al., 2008b). Its design philosophy is to realize high aggregate bandwidth using many moderate-radix switches arranged in a symmetric multi-tier hierarchy, rather than relying on expensive ultra-high-radix switches.
@@
This construction follows the canonical $k$-ary Fat-Tree topology of Al-Fares et al. (Al-Fares et al., 2008a).
@@ 2.5. QFly Architecture
QFly is a flattened, low-diameter switch-centric architecture inspired by Dragonfly-style interconnects, which employ groups of high-radix routers and sparse global links to achieve low diameter while limiting long-cable cost (Kim et al., 2008). In QFly, QPUs are grouped under a smaller number of high-radix optical switches, reducing hierarchy depth and the number of fabric elements compared to Fat-Tree and Clos.
@@ 2.6. BCube Architecture
We follow the original BCube construction of Guo et al. (Guo et al., 2009). An $(n,k_{\mathrm{bcube}})$-BCube consists of $k_{\mathrm{bcube}}+1$ switch layers, each containing $n^{k_{\mathrm{bcube}}}$ switches of radix $n$. Each QPU connects to exactly one switch in every layer, and the total system capacity scales as $n^{k_{\mathrm{bcube}}+1}$ QPUs as shown in figure 5.
@@ 2.7. Optical Loss and EPR-Generation Timing Model
corresponding to a Beneš-style non-blocking optical switch design (Spanke and Benes, 1987). Let $n_{\mathrm{sw}}$ denote the number of switch traversals along the path; then $L_{\mathrm{sw}}=n_{\mathrm{sw}}\,L_{\mathrm{sw}}(k)$.
@@
In server-centric architectures such as BCube, entanglement must be generated on multiple segments and combined through entanglement swapping at intermediate QPUs. Because partially generated entanglement must be stored while waiting for other segments to succeed, we impose a coherence-limited retry window $\tau_{\mathrm{cut}}$. If an end-to-end EPR pair has not been successfully formed via swapping within $\tau_{\mathrm{cut}}$, the attempt is discarded and restarted. We evaluate two repeater-chain scheduling protocols—sequential and parallel—introduced in (Pouryousef et al., 2024), which capture different trade-offs between control simplicity and latency. For these protocols, we estimate the raw end-to-end entanglement delivery rate $R_{\mathrm{raw}}$ and compute
@@ 3. Evaluation
System size is the number of QPUs, denoted by $N$. The total circuit width is set proportional to system scale by assigning each QPU a fixed local grid of data qubits. We evaluate multiple system scales by varying $N$ while keeping the per-QPU data-qubit capacity and the number of communication qubits per QPU fixed. Thus, increasing $N$ increases the total circuit width proportionally (more QPUs, each with the same local qubit grid). The resulting circuits are then partitioned using the same compilation and mapping pipeline and simulated under architecture-specific routing and scheduling policies (we explain this later). Circuits are partitioned using a lightweight Kernighan–Lin (KL) heuristic (Kernighan and Lin, 1970). For the monolithic baseline, we assume a single QPU that hosts all physical qubits required to handle the circuit, supports full all-to-all connectivity, and incurs no inter-QPU communication or entanglement-generation overhead. The circuit execution latency in this case is denoted by $T_{\mathrm{mono}}$.
@@ 3.1. Sanity Checks: Remote-Gate Timing Under Varying Parameters
For the Server-centric paradigm, we evaluate both sequential (S) and parallel (P) swap-scheduling protocols proposed in (Pouryousef et al., 2024), with a memory cutoff time $\tau_{\mathrm{cut}}$ that limits how long a partially generated entangled state may be stored if the required swaps have not yet been completed, after which the attempt is discarded. We report two derived, dimensionless metrics to summarize the behavior of the optical loss and entanglement-timing model.
@@ 3.3. Architecture Instantiations
Following the design space described in Section 2, we instantiate Clos topologies using an automated layout procedure that derives a feasible configuration for each target system size $N$. The procedure selects an even switch radix $k$, determines the number of top-of-rack (ToR) switches $T=k^{2}/4$, and assigns the number of QPUs per rack $R$ such that $T\times R\geq N$, while respecting practical constraints on rack fanout and ToR port budgets. This construction follows the standard folded-Clos model, where each ToR switch allocates $k/2$ ports to QPUs and $k/2$ ports to uplinks (Clos, 1953; Greenberg et al., 2009; Singh et al., 2015).
@@ 3.4. Sensitivity of BCube to communication qubits coherence time
In this experiment, we fix the circuit and network configuration (long-range workload, fixed number of QPUs, communication qubits, and BSM allocation), set the cutoff time $\tau_{\mathrm{cut}}$ in the BCube network model, simulate distributed circuit execution, and record the resulting execution latency. For BCube, we use the parallel swap scheduling protocol proposed in (Pouryousef et al., 2024).